The perimeter of a triangle is 45 feet. The ratio of the width to the length is 2:3:4. Find the measure of each side.
Answer by checkley79(3341)

2+3+4=9 IS THE DIVIDER FOR THE RATIOS.
45/9=5 IS THE MULTIPLIER FOR THE RATIOS.
2*5=10
3*5=15
4*5=20
PROOF:
10+15+20=45
45=45